This Test: 26 pts pr. e. thmthe vite hanno thait they voted in the eas SS% ins, nr. tM2o-ects were asked her wiedinthe election a rind no mean and standard devialen for Po numbers of vmes i, gungs of 1042 b Inthe survey of S142 pe 526 said Deandto㈣decima place as needed) O A. Thi, rend is unusual because S26 is wetm軸range of usual vulues C. This res11b unaual because 526-geat-than tho mainnm usual value O D. This result is unsual because 526 is below the minimum usual vakke O B. Yes.
Answer #1

a) \mu =np =1042*0.55 =573.1

b)\sigma =(np(1-p))1/2 =16.1

b)

as 526 is lower than 2 std deviation from mean:

option D is correct

c)

option A is correct